# Minimax estimation of functionals in sparse vector model with correlated observations

### v1

@@ 6.4.3 Proof of Theorem 8
${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}(\hat{\sigma}^{2}_{\eta}/\sigma^{2}\in(q_{\eta/20}/q_{1-\eta/20},1))\geq 1-\eta/4.$
@@
$\mathcal{A}(\eta):=\left\{\frac{|\hat{\sigma}^{2}_{S}-\sigma^{2}|}{\sigma^{2}}\leq\delta:=c\frac{\|{\boldsymbol{\Sigma}}\|_{F}}{\eta d},\hat{\sigma}^{2}_{\eta}/\sigma^{2}\in(q_{\eta/20}/q_{1-\eta/20},1)\right\}.$
@@ 6.4.4 Proof of Theorem 9
We only prove the lower bound of Theorem 9. The proof of the upper bounds (items (i) and (ii) of Theorem 9) follows from arguments quite analogous to the proofs of Theorems 7 and 8. It is therefore omitted.
Notice that for $s\leq\rho$ the rates stated in the lower bounds of Theorems 4 and 9 are the same while the result of Theorem 4 is stronger since it holds for a smaller minimax risk. This proves the lower bound of Theorem 9 for $s\leq\rho$.
It remains only to prove the lower bound for the case $s>\rho$. Similarly to the proof of Theorem 4, we proceed by a reduction to the lower bound for the model with independent observations. Without causing confusion, we will use the same notation as in Theorem 4 for somewhat different but analogous quantities. Given a fixed loss function $w(\cdot)$ satisfying the assumptions of Theorem 9 and any $\phi>0$, we now define $R^{*}(s,d,\rho,\phi)$ as the minimax risk
$R^{*}(s,d,\rho,\phi):=\inf_{\hat{T}}\sup_{({\boldsymbol{\theta}},{\boldsymbol{\Sigma}})\in\Theta(s,d,\rho)}\sup_{\sigma>0}{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[w\left(\sigma^{-1}\phi^{-1/2}|\hat{T}-\|{\boldsymbol{\theta}}\|_{2}|\right)\right].$
@@
$\bar{\Theta}:=\left\{(\underset{r\;\mathrm{entries}}{\underbrace{\tilde{\theta}_{1},\cdots,\tilde{\theta}_{1}}},\cdots,\underset{r\;\mathrm{entries}}{\underbrace{\tilde{\theta}_{p},\cdots,\tilde{\theta}_{p}}},0,\cdots,0)^{\top}\in\mathbb{R}^{d}:\sum_{i=1}^{p}\mathbbm{1}(\tilde{\theta}_{i}\neq 0)\leq k\right\}.$
By the choice of $p$ we have $d-rp<r$. If the triplet $(r,p,k)$ is such that
@@
$r\phi^{*}(k,p)=\frac{rk}{1\vee\log(k/\sqrt{p})}\gtrsim\frac{s}{1\vee\log(s/\rho)}=\phi^{*}(s,\rho^{2}).$
@@
Let the mixture probability measures $\tilde{\mathbf{P}}_{1},\tilde{\mathbf{P}}_{2}$ and the constant $\tilde{\varphi}>0$ be defined in the same way as $\mathbf{P}_{1}$, $\mathbf{P}_{2}$ and $\varphi$ in the proof of [8, Proposition 3(ii)] with the only difference that $s$ and $d$ are replaced by $k$ and $p$, respectively. We now follow the same argument as in the proof of [8, Proposition 3(ii)] but considering the distributions $\tilde{\mathbf{P}}_{1}\otimes\mathcal{N}(0,1)$ and $\tilde{\mathbf{P}}_{2}\otimes\mathcal{N}(0,1+\tilde{\varphi})$ instead of $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$, respectively. It follows that, in order to prove the lemma, it suffices to show that
@@
$V(\tilde{\mathbf{P}}_{1}\otimes\mathcal{N}(0,1),\tilde{\mathbf{P}}_{2}\otimes\mathcal{N}(0,1+\tilde{\varphi}))$ (180)
$\qquad\leq V(\tilde{\mathbf{P}}_{1}\otimes\mathcal{N}(0,1),\tilde{\mathbf{P}}_{2}\otimes\mathcal{N}(0,1))+V(\tilde{\mathbf{P}}_{2}\otimes\mathcal{N}(0,1),\tilde{\mathbf{P}}_{2}\otimes\mathcal{N}(0,1+\tilde{\varphi}))$ (181)
$\qquad=V(\tilde{\mathbf{P}}_{1},\tilde{\mathbf{P}}_{2})+V(\mathcal{N}(0,1),\mathcal{N}(0,1+\tilde{\varphi})).$ (182)
@@ 6.5 Proof of Proposition 1
As the proof of Theorem 3 we consider two cases. In the first case corresponding to $\sqrt{d}\leq\rho\leq 2\sqrt{d}$ we have $\phi(s,\rho^{2})\leq 2\phi(s,d)$. Then the result follows by reduction to the model with ${\boldsymbol{\Sigma}}={\boldsymbol{I}}$:
${\mathcal{R}}^{*}(s,d,\rho,\bar{r}_{\gamma})\geq\inf_{\Psi}\Big{[}\mathbf{P}_{\mathbf{0},{\boldsymbol{I}}}(\Psi=1)+\sup_{{\boldsymbol{\theta}}\in B_{0}(s,d):\|{\boldsymbol{\theta}}\|_{2}\geq 4\gamma\sigma\sqrt{\phi(s,d)}}\mathbf{P}_{{\boldsymbol{\theta}},{\boldsymbol{I}}}(\Psi=0)\Big{]}\geq c(2\gamma),$ (183)
where $c(\gamma)>0$ is the same constant as in the proof of Theorem 2.
We now focus on the second case $\rho>2\sqrt{d}$. We consider only the sparse regime $s\leq\rho$ since the result in the regime $s>\rho$ follows from an analogous argument. Following the same lines as in the proof of Theorem 3 we find:
@@
${\mathcal{R}}^{*}(s,d,\rho,\bar{r}_{\gamma})\geq\inf_{\Psi}\Big{[}\mathbf{P}_{\mathbf{0},{\boldsymbol{I}}_{p\times p}}(\Psi=1)+\sup_{\tilde{{\boldsymbol{\theta}}}\in B_{0}(1,p):\|\tilde{{\boldsymbol{\theta}}}\|_{2}\geq 2\gamma\sigma\sqrt{c^{\prime}\phi(1,p)}}\mathbf{P}_{\tilde{{\boldsymbol{\theta}}},{\boldsymbol{I}}_{p\times p}}(\Psi=0)\Big{]}\geq c(\sqrt{c^{\prime}}\gamma).$ (186)
Combining the above inequalities and using the definition of $c(\gamma)$ given in the proof of Theorem 2 we obtain the desired result.
Acknowledgment: The work of P. Yang was supported by the grant of National Natural Science Foundation of China (NSFC) 12101353. The work of A.B. Tsybakov was supported by the grant of French National Research Agency (ANR) ”Investissements d’Avenir” LabEx Ecodec/ANR-11-LABX-0047.

### v2

@@ 6.4.3 Proof of Theorem 7
${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}(\hat{\sigma}^{2}_{\eta}/\sigma^{2}\in(q_{\eta/20}/q_{1-\eta/20},1))\geq 1-\eta/4.$ (228)
@@
$\mathcal{A}(\eta):=\left\{\frac{|\hat{\sigma}^{2}_{S}-\sigma^{2}|}{\sigma^{2}}\leq\delta:=c\frac{\|{\boldsymbol{\Sigma}}\|_{F}}{\eta d},\hat{\sigma}^{2}_{\eta}/\sigma^{2}\in(q_{\eta/20}/q_{1-\eta/20},1)\right\}.$ (230)
@@ 6.4.4 Proof of Theorem 8
We only prove the lower bound of Theorem 8. The proof of the upper bounds (items (i) and (ii) of Theorem 8) follows from arguments quite analogous to the proofs of Theorems 6 and 7. It is therefore omitted.
Notice that for $s\leq\rho$ the rates stated in the lower bounds of Theorems 2 and 8 are the same while the result of Theorem 2 is stronger since it holds for a smaller minimax risk. This proves the lower bound of Theorem 8 for $s\leq\rho$.
It remains only to prove the lower bound for the case $s>\rho$. Similarly to the proof of Theorem 2, we proceed by a reduction to the lower bound for the model with independent observations. Without causing confusion, we will use the same notation as in the proof of Theorem 2 for somewhat different but analogous quantities. Given a fixed loss function $w(\cdot)$ satisfying the assumptions of Theorem 8 and any $\phi>0$, we now define $R^{*}(s,d,\rho,\phi)$ as the minimax risk
$R^{*}(s,d,\rho,\phi):=\inf_{\hat{T}}\sup_{({\boldsymbol{\theta}},{\boldsymbol{\Sigma}})\in\Theta(s,d,\rho)}\sup_{\sigma>0}{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[w\left(\sigma^{-1}\phi^{-1/2}|\hat{T}-\|{\boldsymbol{\theta}}\|_{2}|\right)\right].$ (238)
@@
$\bar{\Theta}:=\left\{(\underset{r\;\mathrm{entries}}{\underbrace{\tilde{\theta}_{1},\cdots,\tilde{\theta}_{1}}},\cdots,\underset{r\;\mathrm{entries}}{\underbrace{\tilde{\theta}_{p},\cdots,\tilde{\theta}_{p}}},0,\cdots,0)^{\top}\in\mathbb{R}^{d}:\sum_{i=1}^{p}\mathbbm{1}(\tilde{\theta}_{i}\neq 0)\leq k\right\}.$ (239)
By the choice of $p$ we have drp<rd-rp<r. If the triplet $(r,p,k)$ is such that
@@
$r\phi^{*}(k,p)=\frac{rk}{1\vee\log(k/\sqrt{p})}\gtrsim\frac{s}{1\vee\log(s/\rho)}=\phi^{*}(s,\rho^{2}).$ (248)
@@
[…]
@@
$)\displaystyle V(\tilde{\mathbf{P}}_{1}\otimes\mathcal{N}(0,1),\tilde{\mathbf{P}}_{2}\otimes\mathcal{N}(0,1+\tilde{\varphi}$)) (251)
$)\displaystyle\qquad\leq V(\tilde{\mathbf{P}}_{1}\otimes\mathcal{N}(0,1),\tilde{\mathbf{P}}_{2}\otimes\mathcal{N}(0,1))+V(\tilde{\mathbf{P}}_{2}\otimes\mathcal{N}(0,1),\tilde{\mathbf{P}}_{2}\otimes\mathcal{N}(0,1+\tilde{\varphi}$)) (252)
$).\displaystyle\qquad=V(\tilde{\mathbf{P}}_{1},\tilde{\mathbf{P}}_{2})+V(\mathcal{N}(0,1),\mathcal{N}(0,1+\tilde{\varphi}$)). (253)
@@ 6.5 Proof of Proposition 1
As in the proof of Theorem 2 we consider two cases. In the first case corresponding to $\sqrt{d}\leq\rho\leq 2\sqrt{d}$ we have $\phi(s,\rho^{2})\leq 2\phi(s,d)$. Then the result follows by reduction to the model with ${\boldsymbol{\Sigma}}={\boldsymbol{I}}$ by noticing that
${\mathcal{R}}^{*}(s,d,\rho,\bar{r}_{\gamma})\geq\inf_{\Psi}\Big[\mathbf{P}_{\mathbf{0},{\boldsymbol{I}}}(\Psi=1)+\sup_{{\boldsymbol{\theta}}\in B_{0}(s,d):\|{\boldsymbol{\theta}}\|_{2}\geq 4\gamma\sigma\sqrt{\phi(s,d)}}\mathbf{P}_{{\boldsymbol{\theta}},{\boldsymbol{I}}}(\Psi=0)\Big]\geq{\color[rgb]{0,0,0}c(\gamma)}$ (254)
for all $\gamma>0$, where $c(\gamma)>0$ depends only on $\gamma$. This is done following the same lines as in the proof of [7, Theorem 12]. Note that [7, Theorem 12] obtains (254) for sufficiently small $\gamma$, while we need it here for all $\gamma>0$. Extending the proof in [7] to all $\gamma>0$ is straightforward by using the inequality $V(P,Q)\leq 1-\exp(-\chi^{2}(P,Q))/2$ instead of $V(P,Q)\leq\sqrt{\chi^{2}(P,Q)}$, where $V(P,Q)$ and $\chi^{2}(P,Q)$ are the total variation distance and the chi-square divergence between two probability measures $P$ and $Q$ (cf. [18, Theorem 2.2(iii)]). Following the lines of [7, Theorem 12] with this modification we obtain that the bound (254) holds with
$c(\gamma)=\exp(1-\exp(16\gamma^{2}))/2.$ (255)
We now focus on the second case $\rho>2\sqrt{d}$. We consider only the sparse regime $s\leq\rho$ since the result in the regime $s>\rho$ follows from an analogous argument. Repeating the same steps as in the proof of Theorem 2 we find:
@@
${\mathcal{R}}^{*}(s,d,\rho,\bar{r}_{\gamma})\geq\inf_{\Psi}\Big[\mathbf{P}_{\mathbf{0},{\boldsymbol{I}}_{p\times p}}(\Psi=1)+\sup_{\tilde{{\boldsymbol{\theta}}}\in B_{0}(1,p):\|\tilde{{\boldsymbol{\theta}}}\|_{2}\geq 2\gamma\sigma\sqrt{c^{\prime}\phi(1,p)}}\mathbf{P}_{\tilde{{\boldsymbol{\theta}}},{\boldsymbol{I}}_{p\times p}}(\Psi=0)\Big]\geq{\color[rgb]{0,0,0}c(\sqrt{c^{\prime}}\gamma/2)},$ (258)
where the last inequality follows from the second inequality in (254) that we use with $s$ and $d$ taking values $1$ and $p$, respectively. Combining the above lower bounds for both considered cases and using the expression for $c(\gamma)$ in (255) we obtain the desired result.
Acknowledgment: The work of P. Yang was supported by the grant of National Natural Science Foundation of China (NSFC) 12101353. The work of A.B. Tsybakov was supported by the grants of French National Research Agency (ANR) ”Investissements d’Avenir” LabEx Ecodec/ANR-11-LABX-0047 and ANR MaLip.